Can you help me write the following expression without using the factorial symbol?:
(n-4)!/(n-2)!
thank you

Answer by stanbon(75887)

Can you help me write the following expression without using the factorial symbol?:
(n-4)!/(n-2)!
-----
[(n-4)(n-3)(x-2)!]/(n-2)! = (n-4)(n-3) = n^2 -7n + 12
================================
Cheers,
Stan H.
====================